OpenStudy (anonymous):

Find the simplest factored form of 30x^2(2x+1)^2-10x(2x+1)^3

OpenStudy (anonymous):

Substitute let u =2x+1 and rewrite

OpenStudy (anonymous):

I can't do it like that I have to factor out the LCD

OpenStudy (anonymous):

yes you can \[30x^2u^2-10xu^3\]

OpenStudy (anonymous):

That's easier to see yes? just remember you have to back subsitute

OpenStudy (anonymous):

lcd is \[10xu^ 2 or 10x(2x+1)^2\]

OpenStudy (anonymous):

\[30x^2u^2-10xu^3=10xu^2(3x-u)\] yes?

OpenStudy (anonymous):

then plug 2x+1 back in for u and simplify

OpenStudy (anonymous):

I wish I could do u -sub but it's mandatory that we factor the LCD

OpenStudy (anonymous):

This is what I got

OpenStudy (anonymous):

you are factoring the lcd, youre just calling it U instead of 2x+1

OpenStudy (anonymous):

U is not in the final answer, you have to go back and put in 2x+1 where all the u's are

OpenStudy (anonymous):

10x[3x(2x+1)^2-1(2x+1)^3] =10x(2x+1)^2(3x-1)

OpenStudy (anonymous):

i almost got the same thing i got (x-1) in the final term

OpenStudy (anonymous):

Yeah that's my mistake,

OpenStudy (anonymous):

I don't know what I did wrong?

OpenStudy (anonymous):

let me show you how i did it. watch with U substitution (trust me its easier and its ok to do)

OpenStudy (anonymous):

Okay

OpenStudy (anonymous):

wait I'll try to do it without U substituion but its harder and more typing.

OpenStudy (anonymous):

\[30x^2(2x+1)^2-10x(2x+1)^3=10x(2x+1)^2(3x-2x+1)\]

OpenStudy (anonymous):

oops the last 2x+1 should be in parenthesis, hence -1 instead of +1 in the last term

OpenStudy (anonymous):

Same way but let U = 2x+1 \[30x^2u^2-10xu^3=10xu^2(3x-u)\] back substitute u = 2x+1 \[10x(2x+1)^2(3x-2x-1)=10x(2x+1)^2(x-1)\]

OpenStudy (anonymous):

make sense?

OpenStudy (anonymous):

Yes, thanks

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i might have another problem like that if you want to try one on your own

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

A LOT of problems like this are factored in calculus derivatives, so its good to know how to factor these types of problems. Lets see: ok this one is a bit harder but I think you can do it. Factor \[(x-y)^4-4(x-y)^2\]